Rectangle is 5ft in length more than width the perimeter must be at least 38 feet but no more than 60 feet find the possible dimensions of the width.
Answer by ramkikk66(644)

Rectangle is 5ft in length more than width the perimeter must be at least 38 feet but no more than 60 feet find the possible dimensions of the width.
Let Width be W
Length L = W + 5
Perimeter = 2*(L+W) = 2*(W+W+5) = 4*W + 10
Perimeter must be at least 38 ft. So 4*W + 10 >= 38, or 4*W >= 28
W >= 7 ft. Width must be at least 7 feet
Perimeter must be at most 60 feet. So 4*W + 10 <= 60, or 4*W <= 50
W <= 12.5 ft. Width must be at most 12.5 feet
So possible dimensions of the width are in the range of 7 feet to 12.5 feet.

:)